OpenStudy (anonymous):

A handbag originally cost $85.00. Now it is priced at $68.00. What is the percent of decrease? __ %

OpenStudy (anonymous):

@Nixy

OpenStudy (mrnood):

first work out the decrease in $ i.e. original price - new price post the value here...

OpenStudy (mrnood):

@They_Call_Me_Narii post the difference in price here I can then help you work out the % change

OpenStudy (anonymous):

85.00-68.00=17.00

OpenStudy (anonymous):

Yes, first work out the decrease and we do that by subtracting the the new prices from the original. $85.00 - $68.00 = After getting the difference from above you need to divide the decrease by the original cost and multiply by 100 So we have \( \huge \frac{(85.00 - 68.00)}{85.00} \times 100 = ? \)

OpenStudy (mrnood):

ok so to work out the percentage you need this formula; change in price *100 ____________________ Original price (that is "divided by")

OpenStudy (anonymous):

Also, it will look like this too \( \huge \frac{17}{85.00} \times 100 = ? \)

OpenStudy (anonymous):

\[\frac{ 1700 }{ 8500 }\]